cs161: Data Structures and Algorithms Handout #13 Michael Goldwasser Stanford University Tuesday, July 30, 1996

Homework #5: Dynamic Programming and Introduction to Graphs Due Date: Tuesday, 6 August 1996

Reading

Read Ch. 16, Ch 5.4, Ch. 23.

Practice

Recall, these exercises are purely for your own practice. You need only turn in the official problems.

• Do CLR 16.2-1, 16.2-3.

• Do CLR 16.3-3, 16.3-4.

• Do CLR 16-1. (lengthy)

• Do CLR 23.2-1, 23.2-2.

• Do CLR 23.3-2.

Problems

1. Do CLR 16.3-5.

2. Do CLR 16-5.

3. Do CLR 23.2-5 and CLR 23.3-6. For both parts, you must give a graph as well as a brief desription as to why your graph is a valid solution.

4. Do CLR 23.4-5.

Extra Credit

Do CLR 16.3-6.